The length of a rectangle is three times its width.
If the perimeter of the rectangle is 80ft, find its area.

Answer by JulietG(1812)

You want to put up a perimeter fence around the rectangle. That fencing will require two pieces the length of the land and two pieces the width of the land.
2L + 2W = P
The question gives you the total, 80.
It also gives you the value of the length as 3W.
Let's put in the information:
2(3W) + 2W = 80
Multiply:
6W + 2W = 80
Add:
8W = 80
Divide each side by 8:
W = 10
.
Since we know that the length is three times the W, we can multiply 3*10 to get 30.
The length is 30 and the width is 10.
.
Put this back into the original equation to prove it.
2L + 2W = 80
(2*30)+(2*10) = 80
60 + 20 does indeed equal 80.